OpenStudy (anonymous):

now last question 78x-50=89x+120 value of x find?

OpenStudy (anonymous):

11x=-50-120 11x=-170 x=-170/11

OpenStudy (anonymous):

tnx yar

OpenStudy (anonymous):

78x-50=89x+120 78x -89x=120+50 -11x=170 x=-170/11 x=-15.45

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

:)